A car with a mass of 2500kg and moving at 3 m/s northeast collides with a 3000-kg car moving northwest at 2.5 m/s. find their final velocities. Assume a perfectly inelastic collision.


1
Expert's answer
2021-03-21T11:39:26-0400
"\bold{p_1+p_2=p_f}"

"p_1=(2500)(3)=7500\frac{kgm}{s}\\p_2=(3000)(2.5)=7500\frac{kgm}{s}"

"p_f=(2500+3000)v=2(7500)\sin{45}\\v=1.9\frac{m}{s}"

Direction: north.